OpenStudy (anonymous):

The length of the diagonal of a square is 14. What is the length of each side?

OpenStudy (anonymous):

answers: 7sqrt2 14sqrt2 7sqrt3

OpenStudy (amistre64):

a^2 + b^2 = c^2

OpenStudy (amistre64):

in a square, a = b sooo b^2 + b^2 = c^2 2b^2 = c^2 b^2 = c^2/2 b = c/sqrt(2)

OpenStudy (amistre64):

but it looks like they modified the results such that: c sqrt(2)/2 would suffice

OpenStudy (anonymous):

14 sqrt 2???? Im so confused

OpenStudy (amistre64):

it just the pythagorean thrm

OpenStudy (amistre64):

but since c = 14 14 sqrt(2)/2 would simplify to your answer

OpenStudy (anonymous):

Iknow the pathagorean therum but im not familiar with radicals

OpenStudy (amistre64):

what is 14/2 ?

OpenStudy (anonymous):

7

OpenStudy (amistre64):

then we get 7 sqrt(2) does that match an answer choice?